# Multimodal Evaluation of Russian-language Architectures

### v2

@@ B.16 SchoolScienceVQA
SchoolScienceVQA was developed from scratch based on the methodology of ScienceQA lu2022learn, adapted for Russian cultural and educational context. Domains were adjusted to align with the Russian school curriculum.
@@ B.17 UniScienceVQA
B.18 WEIRD
WEIRD is an extended version of a binary classification subtask of the original English WHOOPS! bittonguetta2023breakingcommonsensewhoops benchmark. The dataset evaluates the ability to detect violations of commonsense. Commonsense violations are situations that contradict the norm of reality DBLP:journals/corr/abs-2503-15948. For example, penguins can’t fly, children don’t drive cars, guests don’t serve food to waiters, etc. “Weird” and “normal” images are equally distributed in the dataset.
@@ B.18 WEIRD
The dataset was created based on the original WHOOPS! bittonguetta2023breakingcommonsensewhoops, using iterative synthetic generation in the style of Self-Instruct rykov-etal-2025-looking. Each sample from the WHOOPS! subset for binary classification is a pair consisting of a “weird” and a “normal” image, along with categories of commonsense violations and image descriptions. To extend the original benchmark, we iteratively generated new categories of commonsense violation and image descriptions using GPT-4o with WHOOPS! samples as a few shots. In addition, we used synthetic descriptions to generate images using DALL-E. Next, we manually filtered out bad images and added good images to the pool. Finally, the pool was used to repeat the generation process and extract new few-shots.
Appendix C Skill taxonomy
Each skill has a modality-related version, while particular skills are not applicable to some modalities, e.g. “Image-to-text grounding” for audio modality. In the description that follows, all skill names are provided along with the corresponding modality tags ( I A V). To disentangle atomic skill names from aggregation taxonomy categories, the former are marked with ➩. Some low-level skills double the names of upper categories, this is to preserve the same taxonomy depth across the entire taxonomy. Descriptions for upper-level skills are inherited by lower levels, unless stated otherwise.
C.1 Perception
Perception involves interpreting and understanding input from different modalities: images, audio, and video. Many perception skills are well-studied in computer vision, speech, and multimodal literature. “Recognition” skills are decomposed as mapping perceived surface features to their corresponding knowledge items, i.e. “naming” objects, events, etc. An overview of the perceprion taxonomy is given in Table 7 and below a detailed skill description is given.
@@ C.1 Perception
Perception
Fine-grained cross-instance perception Overlapping object differentiation Overlapping object differentiation Overlapping image differentiation X
@@
Mutual object localization Mutual object localization Spatial object relationship LabTabVQA X CommonVideoQA
@@
Repeating pattern recognition Repeating pattern recognition Visual pattern recognition X
@@
Event recognition Event recognition Object-object interaction ruCommonVQA ruEnvAQA CommonVideoQA
@@
Fine-grained single-instance perception Object localization Object localization Object localization ruCommonVQA X CommonVideoQA
@@
Object recognition Object recognition Object recognition ruCommonVQA
ruCLEVR ruTiE-Audio CommonVideoQA
RealVQA ruEnvAQA RealVideoQA
ruHHH-Image AQUARIA ruHHH-Video
@@
Event recognition Event recognition Object motion recognition ruCommonVQA ruEnvAQA CommonVideoQA
@@
Textual grounding Image-to-text grounding Image-to-text grounding Scheme recognition ruMathVQA X
@@
Media grounding Media grounding Visual media grounding X
@@
Table 7: Skill taxonomy coverage of MERA Multi tasks (perception part). Columns L1-L5 show the skill hierarchy levels, while Image, Audio and Video columns indicate which tasks cover each knowledge type across different modalities.
C.1.1 Fine-grained single-instance perception
Perceiving and localizing a single object or event.
Object localization
Detecting object positions relative to the image itself (left, right, top, bottom, center). Not to be confused with Media grounding, which deals with direct reference to images and video like frames and bounding boxes.
Object localization
Object localization I V.
Object recognition
@@ C.1.1 Fine-grained single-instance perception
Object recognition
Object recognition I A V.
Single-instance event recognition
Recognizing actions, events, and states. It involves features that are dynamically perceived or may suggest dynamic characteristics, such as those found in motion stop-frames.
Single-instance event recognition
Object motion recognition I A V.
Living things motion recognition I A V.
@@
Non-human pose recognition
Animal body pose recognition I V.
Human pose recognition
Human body pose recognition I V.
Facial expression recognition I V.
Hand gesture recognition I V.
C.1.2 Fine-grained cross-instance perception
@@
Overlapping object differentiation
Disentangling co-occuring objects of the same modality, the ability to separately perceive features of overlapping entitites.
Overlapping image differentiation I V. Separate recognition of visually overlapping objects.
Speaker diarization A V. Attributing speech to multiple simultaneous and/or consequent speakers.
@@
Mutual object localization
Spatial object relationship I A V.
Temporal object relationship A V.
@@
Repeating pattern recognition
Visual pattern recognition I V.
Temporal pattern recognition A V.
@@
Cross-instance event recognition
Object-object interaction I A V.
Human-object interaction I A V.
Human-human interaction I A V.
C.1.3 Textual grounding
@@
Image-to-text grounding
Scheme recognition I V.
Plot recognition I V.
Table recognition I V.
Text recognition (OCR) I V.
@@
Prosody & stress recognition
Prosody & stress recognition A V.
@@
Onomatopoeia A V.
Speech recognition A V.
Song lyrics recognition A V.
@@
Visual media grounding I V.
Temporal media grounding A V.
C.2 Knowledge
C.2.1 Knowledge
@@
Common everyday knowledge I A V.
Ethics I A V.
@@
Common domain knowledge I A V.
Expert domain knowledge I A V.
C.3 Reasoning
@@
Reasoning
Inductive reasoning Attribute recognition Coarse attribute recognition Generated content detection
@@
Object attribute recognition Physical property understanding ruCommonVQA
ruCLEVR
RealVQA ruEnvAQA CommonVideoQA
UniScienceVQA AQUARIA RealVideoQA
@@
Object function understanding ruCommonVQA CommonVideoQA
RealVQA ruEnvAQA RealVideoQA
@@
Other inductive reasoning Other inductive reasoning Other inductive reasoning
Deductive reasoning Deductive reasoning Deductive reasoning Weirdness understanding WEIRD
@@
Abductive reasoning Abductive reasoning Abductive reasoning Hypothetical reasoning RealVQA
@@
Quantitative reasoning Quantitative reasoning Counting Static counting LabTabVQA X CommonVideoQA
@@
Mathematical reasoning Mathematical reasoning ruMathVQA
ruNaturalScienceVQA
RealVQA ruTiE-Audio CommonVideoQA
UniScienceVQA RealVideoQA
@@
Other reasoning Other reasoning Other reasoning Critical thinking
@@
Table 8: Skill taxonomy coverage of MERA Multi tasks (reasoning part). Columns L1-L5 show the skill hierarchy levels, while Image, Audio and Video columns indicate which tasks cover each knowledge type across different modalities.
C.3.1 Inductive reasoning
@@
Generated content detection I A V.
Source characterization I A V.
Media characteristic understanding I A V.
Speech emotion recognition A V.
Music emotion recognition I A V.
Melodic structure interpretation I A V.
Topic understanding I A V.
Style & genre understanding I A V.
Scene understanding I A V.
@@
Physical property understanding I A V.
Object function understanding I A V.
Identity & emotion understanding I A V.
Other inductive reasoning
Other inductive reasoning
Other inductive reasoning I A V.
C.3.2 Deductive reasoning
Deductive reasoning
Deductive reasoning
Weirdness understanding I A V.
Analogical reasoning I A V.
Other deductive reasoning I A V.
C.3.3 Abductive reasoning
Abductive reasoning
Abductive reasoning
Hypothetical reasoning I A V.
Cause & effect understanding I A V.
C.3.4 Quantitative reasoning
Quantitative reasoning

### v3

@@ A.16 SchoolScienceVQA
SchoolScienceVQA was developed from scratch based on the methodology of ScienceQA Lu et al. (2022), adapted for Russian cultural and educational context. Domains were adjusted to align with the Russian school curriculum.
@@ A.17 UniScienceVQA
A.18 WEIRD
WEIRD is an extended version of a binary classification subtask of the original English WHOOPS! Bitton-Guetta et al. (2023) benchmark. The dataset evaluates the ability to detect violations of commonsense. Commonsense violations are situations that contradict the norm of reality Rykov et al. (2025a). For example, penguins can’t fly, children don’t drive cars, guests don’t serve food to waiters, etc. “Weird” and “normal” images are equally distributed in the dataset.
@@ A.18 WEIRD
The dataset was created based on the original WHOOPS! Bitton-Guetta et al. (2023), using iterative synthetic generation in the style of Self-Instruct Rykov et al. (2025b). Each sample from the WHOOPS! subset for binary classification is a pair consisting of a “weird” and a “normal” image, along with categories of commonsense violations and image descriptions. To extend the original benchmark, we iteratively generated new categories of commonsense violation and image descriptions using GPT-4o with WHOOPS! samples as a few shots. In addition, we used synthetic descriptions to generate images using DALL-E. Next, we manually filtered out bad images and added good images to the pool. Finally, the pool was used to repeat the generation process and extract new few-shots.
Appendix B Skill taxonomy
This section provides a detailed description of the skill taxonomy for the MERA Multi benchmark, which was introduced in Section 3.2.
Motivation
This taxonomy is designed to cover the skills required from MLLMs to perform common tasks in multimodal domains.
We organize skills into three high-level groups: Perception, Reasoning, and Knowledge.
  • Perception covers extraction of salient information from images/audio/video;
  • Reasoning covers inference over extracted information (often combining multiple cues and background assumptions);
  • Knowledge covers retrieval and application of stored factual or domain information.
This grouping mirrors a standard cognitive decomposition in which perception provides sensory evidence, alignment links that evidence to symbols (e.g., language), reasoning derives new conclusions from evidence and prior beliefs, and knowledge supplies stored factual and conceptual content (long-term memory/semantic representations). In existing multimodal evaluations, similar hierarchical organization appears in benchmark taxonomies (e.g., ConvBench’s perception→reasoning→creativity hierarchy Liu et al. (2024d); MMBench’s fine-grained perception categories Liu et al. (2024e); MMStar’s perception-reasoning-knowledge hierarchy Chen et al. (2024); MME’s perception-cognition Fu et al. (2024)), while knowledge-heavy competence is frequently assessed through QA/exam-style benchmarks such as MMLU/MMMU and cognitively motivated “core knowledge” suites such as CoreCognition Li et al. (2025).
Structure
The core skill set is in the last level of the taxonomy, which is called atomic skills. Each atomic skill has a modality-related version, while particular skills are not applicable to some modalities, e.g. “Image-to-text grounding” is irrelevant for audio modality. In the description that follows, all skill names are provided along with the applicable modality tags ( I A V). To disentangle atomic skill names from aggregation taxonomy categories, the former are marked with ➩ in the following section.
Taxonomy tables (Table 3, Table 7 and Table 8) provide a structured overview of the skill taxonomy along with their coverage by MERA Multi tasks. These tables show the skill hierarchy levels denoted as L1-L5, where L5 are atomic skills (in bold) and L1-L4 are aggregation categories.
@@ Appendix B Skill taxonomy
Perception Fine-grained cross-instance perception Overlapping object differentiation Overlapping image differentiation X
@@
Mutual object localization Spatial object relationship LabTabVQA X CommonVideoQA
@@
Repeating pattern recognition Visual pattern recognition X
@@
Event recognition Object-object interaction ruCommonVQA ruEnvAQA CommonVideoQA
@@
Fine-grained single-instance perception Object recognition Object localization ruCommonVQA X CommonVideoQA
@@
Object recognition ruCommonVQA ruTiE-Audio CommonVideoQA
ruCLEVR ruEnvAQA RealVideoQA
RealVQA AQUARIA ruHHH-Video
ruHHH-Image
@@
Event recognition Object motion recognition ruCommonVQA ruEnvAQA CommonVideoQA
@@
Textual grounding Image-to-text grounding Scheme recognition ruMathVQA X
@@
Media grounding Visual media grounding X
@@
Table 7: Skill taxonomy coverage of MERA Multi tasks (perception part). Columns L1-L5 show the skill hierarchy levels, while Image, Audio and Video columns indicate which tasks cover each perception type across different modalities.
B.1 Perception
Perception denotes the set of abilities by which a model extracts task-relevant information from sensory inputs (image, audio, video) and converts it into internally usable representations. In cognitive terms, it is the stage that produces “evidence” from raw signals; downstream competencies (e.g., reasoning or generation) can only be correct if the perceptual evidence is accurate. This dependency is made explicit in hierarchical benchmark taxonomies such as ConvBench Liu et al. (2024d) (perception→reasoning→creativity) and MMStar Chen et al. (2024). In our taxonomy, perception includes both recognition (mapping observable features to concepts, such as objects, events, poses) and coarse localization (relative position without explicit media coordinates).
We define 3 broad categories of perception skills. The dichotomy between single-instance and cross-instance perception has been popularized in MMBench Liu et al. (2024e) and subsequently reused by HR-Bench Wang et al. (2024b). The third category is multimodal alignment which includes skills for matching multimodal representations, including text recognition.
An overview of the perception taxonomy is given in Table 7 and below a detailed skill description is given.
B.1.1 Fine-grained single-instance perception
Perceiving and (optionally) localizing a single salient object/event instance in an input.
Object localization I V
Detecting object positions relative to the image itself (left, right, top, bottom, center) and predicts coarse spatial categories (e.g., left/right/center) without producing media coordinates. In contrast, media grounding predicts locations in the media coordinate system (bounding boxes, segmentation masks, or timestamp spans).
Object recognition I A V
@@ B.1.1 Fine-grained single-instance perception
Event recognition (single-instance)
Recognizing actions, events, and states. It involves features that are dynamically perceived (e.g. rolling) or may suggest dynamic characteristics (e.g. color changes), such as those found in motion stop-frames.
Object motion recognition I A V
Living things motion recognition I A V
Recognizing movement types specific to humans and animals.
@@
The ability to infer articulated body or hand configuration, either as keypoints/joint structure or as gesture/pose classes when the task is categorical. It deserves a separate perceptual skill because many downstream tasks (e.g. action and intention understanding) depend primarily on how a person is posed rather than what objects are present. In practice, this is operationalized by standard pose-estimation datasets such as COCO Keypoints Lin et al. (2014) and MPII Human Pose Andriluka et al. (2014), video benchmarks like PoseTrack Andriluka et al. (2018) for multi-person pose and tracking, and hand-gesture datasets such as HaGRID Kapitanov et al. (2024), for detection- and classification-style gesture recognition.
Body pose recognition I V
Facial expression recognition I V
Hand gesture recognition I V
B.1.2 Fine-grained cross-instance perception
@@
Separating co-occurring entities whose features interfere.
Overlapping image differentiation I V
Disentangling and separate recognition of visually overlapping objects, .
Speaker diarization A V
Attributing speech to the correct speaker among multiple simultaneous and/or consequent speakers.
@@
Perceiving relative relations among multiple entities.
Spatial object relationship I A V
Understanding the relative localization of multiple objects in the scene.
Temporal object relationship A V
Understanding the ordering and temporal relations across events/segments.
@@
Perceiving instances that appear as a repeating pattern, e.g. animal flocks, ornaments, rhythmic sounds and moves, also involves perceiving frequency and density. Primary to counting skills.
Visual pattern recognition I V
Patterns present in a single image or video frame, which do not require dynamic perception to acquire.
Temporal pattern recognition A V
Patterns spanning over time and / or perceived only across multiple frames.
@@
Perceiving interactions among multiple entities.
Object-object interaction I A V
Human-object interaction I A V
Human-human interaction I A V
B.1.3 Cross-modal alignment
Cross-modal alignment (Grounding) covers abilities that establish correspondences between symbolic descriptions (typically language) and specific elements in a non-text modality, enabling verification (“is the described thing present?”) and reference (“where/when is it?”). In cognitive terms, grounding mediates between perceptual evidence and symbolic reasoning, supporting tasks that require referential linkage rather than mere recognition. We separate two complementary families: (i) Text extraction — recovering textual symbols from media (e.g., OCR in images/video frames; speech recognition and lyrics recognition in audio/video) — and (ii) Media grounding — mapping language to media coordinates (e.g., bounding boxes/masks in images/video, or temporal spans in audio/video). This separation keeps transcription distinct from localization while preserving a unified notion of cross-modal correspondence.
@@ B.1.3 Cross-modal alignment
Involves recognizing text in images, video frames, and more complex visual elements such as tables, charts, and diagrams. Some of the benchmarks targeting these skills together with QA tasks are PlotQA Methani et al. (2020), ChartQA Masry et al. (2022), Misleading ChartQA Chen et al. (2025), ChartBench Xu et al. (2023), and TextVQA Singh et al. (2019).
Text recognition (OCR) I V
Scheme recognition I V
Plot recognition I V
Table recognition I V
@@
Similarly to Image-to-text grounding, this skill involves matching sound signals to textual representation, including complex audio elements such as prosody, stress, and lyrics.
Prosody & stress recognition A V
involves recognizing and understanding the rhythm, intonation, and stress patterns in speech.
@@
Onomatopoeia A V
Onomatopoeia is a language phenomenon where words are used to imitate sounds, such as “buzz” for a bee or “meow” for a cat. Recognizing onomatopoeia is essential for understanding routine language usage. An example of a dataset for onomatopoeia recognition is RWCP-SSD-Onomatopoeia Okamoto et al. (2020).
Speech recognition A V
Song lyrics recognition A V
@@
Grounding to media involves matching textual descriptions to media elements via masking, segmentation, or bounding boxes.
Visual media grounding I V
Operating on bounding boxes, segmentation masks, or pixel coordinates for objects or actions described in natural language. For example, predicting the bounding box of the object in an image or the segmentation mask of the action in a video.
Temporal media grounding A V
Operating on timestamps and media duration. For example, predicting the start and end time of the event in a video or the duration of the action in a video.
B.2 Knowledge
Knowledge denotes a model’s ability to retrieve, apply, and integrate stored factual and conceptual information in multimodal settings. In cognitive science terms, this corresponds to long-term semantic memory (stable representations of entities, relations, norms, and domain concepts) used to interpret perceptual evidence and to support reasoning. In multimodal evaluation, knowledge competence is often probed via QA and exam-style benchmarks (e.g., MMLU/MMMU hendrycks2021measuring; Yue et al. (2024)) that condition questions on images, and via cognitively motivated suites that target foundational concepts from developmental cognition (e.g., CoreCognition Li et al. (2025), which operationalizes “core knowledge” constructs such as object permanence and basic physical/social regularities). Because “knowledge” interacts with both perception and reasoning, we treat it as a distinct group to separate failures of recall/concept access from failures of inference over correctly perceived evidence.
@@ B.2 Knowledge
In MLLM benchmarks, “knowledge” is typically observed as (i) reliance on commonsense/everyday facts in visual contexts, (ii) explicit external/encyclopedic knowledge conditioned on images, and (iii) expert-domain question answering over multimodal inputs.
The differentiation between common and domain knowledge may vary depending on the task and the domain. We formulate the polar points of this distinction as follows: common knowledge is generally agreed upon by the majority of the population and used in everyday life, while expert knowledge is domain-specific and requires specialized knowledge or training.
@@
Common everyday knowledge I A V
Commonsense and broadly shared facts.
Ethics I A V
Safety/ethics-aligned judgment about harmful/unsafe content or behavior in multimodal prompts. The placement of ethics in the knowledge category and treating ethical norms as stored conventions may be considered a simplification, but it is done to keep minimal the number of major categories.
@@
This category may be further subdivided into subject-specific categories as it is done in expert benchmarks, e.g. MMMU Yue et al. (2024).
Common domain knowledge I A V
Subject-specific but broadly accessible knowledge (often school-level or general STEM).
Expert domain knowledge I A V
Specialized, exam- or professional-level knowledge requiring training.
B.3 Reasoning
Reasoning comprises abilities that derive new conclusions from (i) perceptual evidence and (ii) prior knowledge, including uncertain, multi-step, relational, causal, quantitative, and counterfactual inference. In cognitive terms, reasoning corresponds to controlled inference processes that operate on representations supplied by perception and memory, producing answers that are not directly observable in the input. We include both classical forms (deductive/abductive/inductive inference) and task-driven “holistic” judgments (e.g., topic, style, provenance, media characteristics) when they require integrating multiple cues across an input rather than naming a single localized entity. This placement is consistent with benchmark taxonomies that separate fine-grained recognition from higher-level judgments—for example, MMBench distinguishes coarse perception categories such as image topic/quality from fine-grained object-level recognition Liu et al. (2024e). Analogously, audio evaluation benchmarks emphasize media-level characterization and instruction-guided understanding (AudioBench Wang et al. (2025a), AIR-Bench Yang et al. (2024), MMAU Sakshi et al. (2025)).
@@ B.3 Reasoning
Reasoning Inductive reasoning Attribute recognition Coarse attribute recognition Generated content detection
@@
Object attribute recognition Physical property understanding ruCommonVQA ruEnvAQA CommonVideoQA
ruCLEVR AQUARIA RealVideoQA
RealVQA
UniScienceVQA
@@
Object function understanding ruCommonVQA ruEnvAQA CommonVideoQA
RealVQA RealVideoQA
@@
Other inductive reasoning Other inductive reasoning
Deductive reasoning Weirdness understanding WEIRD
@@
Abductive reasoning Hypothetical reasoning RealVQA AQUARIA CommonVideoQA
@@
Quantitative reasoning Counting Static counting LabTabVQA X CommonVideoQA
@@
Mathematical reasoning Mathematical reasoning ruMathVQA ruTiE-Audio CommonVideoQA
ruNaturalScienceVQA RealVideoQA
RealVQA
UniScienceVQA
@@
Other reasoning Critical thinking
@@
Table 8: Skill taxonomy coverage of MERA Multi tasks (reasoning part). Columns L1-L5 show the skill hierarchy levels, while Image, Audio and Video columns indicate which tasks cover each reasoning type across different modalities.
B.3.1 Inductive reasoning
@@
The distinction between coarse attribute recognition and object attribute recognition is inspired by MMBench’s Liu et al. (2024e) categorization into coarse vs. fine-grained perception. We put coarse perception into reasoning category as most of the related tasks require inference from the perceived information about the whole sample.
@@
Global or “holistic” characterization of a sample (quality, topic, provenance, modality-level traits), without requiring fine object-level localization or multi-entity comparison.
Topic understanding I A V
Style & genre understanding I A V
Scene understanding I A V
Generated content detection I A V
Detection of whether content is synthetic vs real.
Media characteristic understanding I A V
Recognize broad media-level traits (quality, augmentations). Characterizing/attributing the origin or type of synthesis (e.g., illustration type, camera quality, sound interference).
In vision-language evaluation, MMBench Liu et al. (2024e) explicitly includes Image Quality and Image Topic as coarse perception categories. In audio(-text) evaluation, media-level characterization is a core target of AudioBench Wang et al. (2025a) (speech understanding, audio scene understanding, and paralinguistic “voice understanding”), AIR-Bench Yang et al. (2024) (speech/natural sounds/music comprehension under instruction following), and MMAU Sakshi et al. (2025) (multi-task audio understanding and reasoning across speech/environment/music).
Speech emotion recognition A V
Music emotion recognition I A V
Melodic structure interpretation I A V
@@
Attribute-centric understanding grounded in entities (objects, people, instruments), including physical properties, functions, and identity-related cues.
Physical property understanding I A V
Infer physical properties (e.g., material/shape/size/texture; or physical dynamics in video).
Object function understanding I A V
Infer what an object is for and how it is used (affordances, intent).
Identity & emotion understanding I A V
Recognize identity-related cues (who) and affective state (facial expression, vocal affect).
B.3.2 Deductive reasoning
Rule- or structure-driven inference where conclusions follow from constraints (logical, relational, symbolic, or structural).
Weirdness understanding I A V
Detect incongruity, humor, anomalies, or “oddness” that violates expectations.
Analogical reasoning I A V
Solve problems by mapping relational structure between situations.
Other deductive reasoning I A V
General logical/relational deduction (multi-hop, constraints).
B.3.3 Abductive reasoning
Inference to the best explanation: forming hypotheses and causal attributions consistent with observations.
Hypothetical reasoning I A V
Reason about “what if” interventions and unseen outcomes.
Cause & effect understanding I A V
Identify causal responsibility and predict effects of interactions over time.
B.3.4 Quantitative reasoning